OpenStudy (anonymous):

Simplify completely quantity 2 x squared plus 13 x plus 20 all over x squared minus 5 x minus 36

OpenStudy (anonymous):

@ranga

OpenStudy (anonymous):

@tkhunny

OpenStudy (tkhunny):

Have you considered factoring? \(\dfrac{2x^{2}+13x+20}{x^{2}-5x-36} = \dfrac{(x+4)(2x+5)}{(x+4)(x-9)}\) I see a clue in there.